# Black-hole thermodynamics in doubly special relativity: universal g/f Hawking-temperature scaling

### v2

@@ Abstract
Doubly Special Relativity (DSR) deforms special-relativistic kinematics while preserving a relativity principle by introducing, in addition to the invariant speed of light, a second invariant scale that is typically identified with the Planck energy $E_{\mathrm{Pl}}$. In applications to curved spacetimes and, in particular, to black-hole thermodynamics, a persistent conceptual and operational issue concerns the meaning of the “energy” entering DSR-inspired modified dispersion relations (MDRs): different identifications (Killing energy at infinity, energy measured by a local observer, or a phase-space invariant) can lead to superficially different predictions and may effectively reintroduce preferred frames if implemented inconsistently.
In a controlled static black-hole setting we compare two widely used curved-spacetime implementations: (i) MDRs imposed in local orthonormal frames on an energy-independent background geometry, and (ii) the rainbow-metric proposal in which the deformation is encoded in an energy-dependent family of effective metrics. Restricting to static, spherically symmetric horizons and adopting a consistent finite operational scale $E_{\star}$ associated with the emitted quanta, we show that both prescriptions yield the same near-horizon temperature rescaling,
$T(E_{\star})=T_{0}\,\frac{g(E_{\star}/E_{\mathrm{Pl}})}{f(E_{\star}/E_{\mathrm{Pl}})}\,,\qquad T_{0}=\frac{\kappa_{0}}{2\pi}\,$, (1)
where $f$ and $g$ are the standard rainbow/MDR functions. The result can be viewed as a universality statement for the tunneling/surface-gravity temperature: within these assumptions, the deformation enters only through the ratio $g/f$.
We illustrate the scaling for an Amelino–Camelia-type MDR and for the original Magueijo–Smolin DSR invariant, for which $f=g$ and therefore $T(E_{\star})=T_{0}$ at this level. We then extend the analysis to a two-parameter generalized DSR family (G-DSR/GDRS) characterized at leading order by $(\alpha_{2},\Delta\alpha)$, obtaining
@@
We discuss the interpretation of the controlling combination $\Delta\alpha-\alpha_{2}$, the special role of the symmetric subfamily $\Delta\alpha=\alpha_{2}$ (for which the leading temperature correction vanishes), and the sources of model dependence that lie beyond the near-horizon temperature—notably phase-space measures, greybody factors, massive thresholds, and (in fully fledged DSR) non-linear energy-composition laws. For macroscopic black holes, any correction associated with $E_{\star}/E_{\mathrm{Pl}}$ remains strongly suppressed; consequently, quantitative effects are expected to become significant only near the Planck regime, where backreaction and genuinely quantum-gravitational dynamics should determine the endpoint.
@@ 1 Introduction
Hawking’s discovery that black holes radiate [1], together with the Bekenstein area law for entropy [2], established a thermodynamic interpretation of horizon mechanics in which the Hawking temperature is set by the surface gravity, $T_{0}=\kappa_{0}/(2\pi)$ (in units $G=c=\hbar=k_{B}=1$). Comprehensive treatments may be found in [3, 4]. Black-hole thermodynamics thus provides a uniquely constrained arena for probing how semiclassical quantum field theory interfaces with gravity and for assessing the robustness of universal horizon phenomena under ultraviolet (UV) modifications.
@@
A long-standing motivation for considering Planck-scale modifications is the so-called trans-Planckian problem: in standard derivations, outgoing Hawking quanta originate from modes that, when traced back towards the horizon, attain arbitrarily large frequencies in the local static frame. A substantial body of work indicates that the late-time Hawking flux is remarkably insensitive to many UV modifications of the microphysics, including broad classes of modified dispersion relations and analogue-gravity models [5, 6, 7, 8, 9]. Nevertheless, UV deformations remain useful as effective parameterizations of departures from standard local quantum field theory, and they may influence not only the inferred near-horizon temperature but also the detailed spectrum through modified group velocities, densities of states, and scattering (greybody) factors.
In addition to the original Bogoliubov-coefficient approach, several complementary derivations connect the temperature to near-horizon analyticity and to the geometry of the horizon. Particularly relevant for our purposes is the tunneling picture, in both the null-geodesic and Hamilton–Jacobi formulations [10, 11, 12]. These methods make explicit how the pole structure of the radial momentum near the horizon encodes the thermal factor and are well suited for incorporating MDRs at the level of a Hamiltonian constraint.
Doubly Special Relativity (DSR) provides a well-developed framework for deforming kinematics while preserving a relativity principle by introducing a second invariant scale, typically the Planck energy $E_{\mathrm{Pl}}$ [13, 14, 15, 16, 17]. In DSR, modified dispersion relations arise alongside non-linear Lorentz transformations and, in general, deformed multi-particle composition laws. A key virtue of DSR, as opposed to generic Lorentz-violating effective field theories, is that it aims to avoid a preferred frame; however, this virtue becomes subtle in curved spacetime, where the operational meaning of the energy scale entering the deformation must be specified with care.
The curved-spacetime ambiguity.
The central difficulty for black-hole applications is that the energy entering the MDR is not uniquely defined in a curved background. In a stationary spacetime one can use the conserved Killing energy $E_{\infty}$, but local energies measured by static observers diverge at the horizon. Alternatively, one may use energies measured in freely falling orthonormal frames that remain regular across the horizon. The choice is not merely a matter of notation: because DSR deformations are written as functions of $E/E_{\mathrm{Pl}}$, different operational identifications can lead to different numerical predictions and, if handled inconsistently, can effectively reintroduce a preferred frame.
Implementations compared in this paper.
Two prescriptions dominate much of the DSR/rainbow black-hole literature:
  • Fixed background with local-frame MDR: the spacetime geometry is classical and energy independent, while particle propagation is governed by an MDR imposed in a local orthonormal frame.
  • Rainbow gravity: the deformation is encoded in an energy-dependent family of effective metrics labelled by an energy scale associated with the probe [18].
While often presented as conceptually distinct, these prescriptions become closely related in static, near-horizon computations once the energy scale entering the deformation is specified operationally and implemented consistently.
Objective and main message.
The objective of this paper is twofold. First, we provide an explicit comparison of fixed-background (local-frame) MDR computations and rainbow-metric computations of the Hawking temperature for static, spherically symmetric horizons. Second, we show that, under a common and physically motivated assumption that the MDR/rainbow functions are evaluated at the same finite operational scale $E_{\star}$, both prescriptions yield the same temperature rescaling through the universal ratio $g/f$. In this sense, the primary source of apparent discrepancies across the literature is not the choice between “local MDR” and “rainbow metric” parametrizations, but rather the prescription used to identify $E_{\star}$.
Generalized DSR (G-DSR/GDRS).
@@
Beyond the standard Amelino–Camelia (AC) and Magueijo–Smolin (MS) realizations, we incorporate a two-parameter generalized DSR family (G-DSR/GDRS) with leading-order parameters $(\alpha_{2},\Delta\alpha)$ [30]. This parametrization makes transparent which combinations of deformation parameters control (or eliminate) the leading-order Hawking-temperature correction in the $g/f$ scaling.
Organization.
@@
Section 2 reviews MDR conventions and representative DSR realizations. Section 3 discusses curved-spacetime extensions and highlights the operational energy issue. Section 4 derives the universal $g/f$ temperature scaling for static horizons using both local-frame MDR and rainbow-metric viewpoints, emphasizing their equivalence under a common scale choice. Section 5 applies the result to AC, MS, and G-DSR/GDRS and discusses simple thermodynamic implications in the Schwarzschild case. Section 6 addresses sources of model dependence beyond the near-horizon temperature, and Section 7 provides a focused discussion of the generalized DSR extension and the interpretation of results. We conclude with a summary and outlook.
@@ 2.3 Magueijo–Smolin (MS) DSR invariant
so that $g_{\rm MS}/f_{\rm MS}=1$. This equality will play an important role below: even though the MDR is non-trivial, the near-horizon temperature correction vanishes in the universal $g/f$ scaling.
@@ 3.4 The central ambiguity: which energy enters $E_{\star}/E_{\mathrm{Pl}}$?
4 Static horizons and the universal $g/f$ scaling of the Hawking temperature
@@ 6 Beyond the near-horizon temperature: sources of model dependence
The universality of (23) at the surface-gravity/tunneling level does not imply that all DSR-motivated models lead to identical evaporation phenomenology. The Hawking temperature captures the near-horizon analyticity and redshift structure, but the observable flux at infinity depends on additional inputs. For a generic species $i$ with greybody factor $\Gamma_{i}(\omega)$, the standard energy flux has the schematic form
@@
Sources of model dependence.
Important effects include:
  • Massive thresholds and kinematics: emission thresholds and the relation between energy, momentum, and group velocity can modify fluxes even if $T(E_{\star})$ is unchanged (e.g. when $f=g$).
  • Phase-space measure and density of states: in DSR, curved momentum space can induce a non-trivial integration measure, affecting densities of states and spectral distributions [26, 24].
  • Greybody factors: the scattering problem in the exterior geometry may be altered if the effective wave equation and group velocity are modified; this changes $\Gamma_{i}(\omega)$ and can dominate over modest temperature rescalings.
  • Energy-composition laws and backreaction: in fully fledged DSR, multi-particle kinematics involves non-linear composition laws, which can influence how emitted quanta contribute to energy balance and, potentially, to backreaction and energy conservation in evaporation scenarios.
These considerations clarify why agreement at the level of the near-horizon temperature does not guarantee agreement for integrated evaporation histories or for endpoint behavior.
Operational scale revisited.
Even within the temperature calculation, the choice of $E_{\star}$ is the dominant ambiguity. If $E_{\star}$ is identified with a detector energy at infinity, then the deformation may effectively describe a frequency-dependent reweighting of the observed spectrum. If $E_{\star}$ is tied self-consistently to the local emission process (e.g. to $T$ itself), the deformation becomes mass-dependent and can qualitatively change the late stages of evaporation in models where $g/f$ approaches zero at finite argument [21, 20]. A meaningful comparison across approaches therefore requires a complete specification of the operational prescription for $E_{\star}$ and of the dynamical and statistical ingredients beyond the near-horizon temperature.
@@ 7 Discussion: implications of the G-DSR/GDRS extension and interpretation of results
Interpretation of the universal $g/f$ scaling.
Our main technical result can be interpreted as follows. For static horizons, the temperature is governed by two ingredients: (i) the geometric pole structure in the radial momentum near the horizon, and (ii) the relation between energy and radial momentum implied by the MDR. Under the operational assumption that the deformation functions are evaluated at a fixed finite scale $E_{\star}$, ingredient (ii) enters only as an overall rescaling. Rainbow metrics implement precisely the same rescaling through an energy-dependent normalization of time and space in the orthonormal frame. Therefore, at the level of near-horizon temperature, the two prescriptions are equivalent parametrizations.
Scope and limitations.
@@ 8 Conclusion
We have compared two standard prescriptions for implementing DSR-motivated modified dispersion relations in black-hole spacetimes: (A) an energy-independent background with the MDR imposed in local orthonormal frames, and (B) the rainbow-metric construction in which the deformation is encoded in an explicitly energy-dependent effective geometry. For static, spherically symmetric horizons, we showed explicitly that both prescriptions lead to the same modification of the Hawking temperature at the surface-gravity/tunneling level when expressed in the ffgg parametrization and evaluated at the same finite operational scale $E_{\star}$,
@@
At this level, the distinction between “local-frame MDR on a fixed background” and “rainbow metric” is primarily a distinction of parametrization, provided the physical prescription for $E_{\star}$ is held fixed.
@@
Finally, we emphasized that the near-horizon temperature is only one component of the evaporation problem. Threshold effects, phase-space measures, greybody factors, and (in full DSR) deformed composition laws can generate substantial model dependence even when $T(E_{\star})$ coincides. In practice, the dominant driver of quantitative differences across approaches is the operational identification of the deformation scale $E_{\star}$ and the corresponding physical interpretation of the “energy” entering the deformation functions. For macroscopic black holes, corrections remain strongly suppressed by $E_{\star}/E_{\mathrm{Pl}}$ and are expected to become relevant only near the Planck regime, where backreaction and full quantum-gravity dynamics should control the endpoint.

### v3

# Black-hole thermodynamics in doubly special relativity: near-horizon $g/f$ temperature scaling under a shared operational scale

@@ Abstract
Doubly Special Relativity (DSR) deforms special-relativistic kinematics while preserving a relativity principle by introducing, in addition to the invariant speed of light, a second invariant scale that is typically identified with the Planck energy $E_{\mathrm{Pl}}$. In curved spacetime, however, the physical meaning of the “energy” entering DSR-inspired modified dispersion relations (MDRs) is ambiguous: Killing energies, local observer energies, and phase-space invariants need not coincide. Consequently, for black-hole applications, this ambiguity can be just as important as the algebraic form of the MDR itself.
The main objective of this paper is to clarify, within a single and consistent framework, whether two widely used curved-spacetime implementations of DSR deformations really lead to different Hawking-temperature predictions or merely represent different parametrizations of the same near-horizon effect. To this end, we compare (i) MDRs imposed in local orthonormal frames on an energy-independent background geometry and (ii) the rainbow-metric proposal, in which the deformation is encoded in an energy-dependent family of effective metrics. Restricting to static, spherically symmetric horizons and, moreover, evaluating both constructions at the same finite operational scale $E_{\star}$, we show that the two prescriptions lead to the same near-horizon temperature rescaling,
Therefore, our result is not a first-principles derivation of a universal DSR prediction; rather, it is a controlled equivalence statement showing that, under a shared operational prescription, the local-MDR and rainbow descriptions amount to the same parametrization of the near-horizon temperature shift.
We then illustrate this statement for an Amelino–Camelia-type MDR and for the original Magueijo–Smolin DSR invariant, for which $f=g$ and therefore $T(E_{\star})=T_{0}$ at this level. Next, we extend the comparison to a two-parameter generalized DSR family (G-DSR/GDRS) characterized at leading order by $(\alpha_{2},\Delta\alpha)$, obtaining
@@
In this way, we make explicit that the leading correction is controlled by the single combination $\Delta\alpha-\alpha_{2}$, which vanishes on the symmetric subfamily $\Delta\alpha=\alpha_{2}$. Finally, we show how the size of the correction depends on the prescription used for $E_{\star}$: for the common choice $E_{\star}\sim T_{0}$, the relative shift is suppressed by $1/(ME_{\mathrm{Pl}})$ and is therefore negligible for macroscopic black holes. Accordingly, observable or phenomenological consequences must come either from Planckian regimes or from ingredients beyond the near-horizon temperature, such as greybody factors, phase-space measures, thresholds, and deformed multi-particle composition laws.
@@ 1 Introduction
Hawking’s discovery that black holes radiate [1], together with the Bekenstein area law for entropy [2], established a thermodynamic interpretation of horizon mechanics in which the Hawking temperature is set by the surface gravity, $T_{0}=\kappa_{0}/(2\pi)$ (in units $G=c=\hbar=k_{B}=1$). Comprehensive treatments may be found in [3, 4]. Black-hole thermodynamics thus provides a constrained setting for testing how semiclassical quantum field theory responds to ultraviolet (UV) modifications.
@@
A standard motivation for Planck-scale deformations is the trans-Planckian problem: in conventional Hawking calculations, outgoing quanta originate from modes that acquire arbitrarily large local frequencies when traced backwards toward the horizon. A large body of work nevertheless suggests that the late-time Hawking flux is rather robust under many UV deformations of the microphysics, including broad classes of modified dispersion relations and analogue-gravity constructions [5, 6, 7, 8, 9]. Even so, MDR-based effective descriptions remain useful because they can modify the inferred near-horizon temperature, the propagation of wave packets, and the detailed spectrum through group velocities, densities of states, and greybody factors.
Doubly Special Relativity (DSR) is an especially interesting framework in this context because it aims to deform relativistic kinematics while preserving a relativity principle through the introduction of a second invariant scale, typically the Planck energy $E_{\mathrm{Pl}}$ [13, 14, 15, 16, 17]. In flat spacetime, MDRs are then accompanied by non-linear Lorentz transformations and, in general, deformed composition laws for multi-particle systems. In curved spacetime, however, one immediately faces a serious operational problem: the energy entering the deformation is not uniquely defined. One may use a conserved Killing energy at infinity, a local energy measured by static or freely falling observers, or a phase-space invariant. Different choices need not agree numerically, and if they are mixed inconsistently they can effectively reintroduce a preferred frame.
This ambiguity is central to the comparison between the two curved-spacetime prescriptions that dominate much of the DSR/rainbow black-hole literature. In the first, the spacetime geometry is kept classical and energy independent, while particle propagation is governed by an MDR imposed in a local orthonormal frame. In the second, the deformation is encoded in an energy-dependent family of effective metrics, as in rainbow gravity [18]. Against this background, the main objective of the present work is to determine whether these two procedures genuinely yield different Hawking-temperature predictions or, instead, become equivalent once the same physical energy prescription is adopted. More precisely, we show that for static, spherically symmetric horizons, and provided that the same finite operational scale $E_{\star}$ is used in both constructions, they lead to the same near-horizon temperature rescaling by the ratio $g/f$.
We emphasize from the outset that this is a conditional equivalence statement, not a derivation of a unique DSR prediction from first principles. Rather, what is being shown is that once a common prescription for the deformation scale is imposed, the apparent difference between “local MDR” and “rainbow metric” descriptions largely reduces to parametrization at the level of the tunneling/surface-gravity temperature. Therefore, this framing is important because the operational definition of $E_{\star}$ remains the dominant source of ambiguity in curved-spacetime DSR applications.
@@
Besides the standard Amelino–Camelia (AC) and Magueijo–Smolin (MS) realizations, we also incorporate a two-parameter generalized DSR family (G-DSR/GDRS) with leading-order parameters $(\alpha_{2},\Delta\alpha)$ [30]. In this way, the analysis makes transparent which combination of deformation parameters controls the leading correction and which subfamily leaves the near-horizon temperature unchanged.
@@
The paper is organized as follows. Section 2 reviews the MDR conventions and the representative DSR realizations used in the comparison. Section 3 discusses curved-spacetime extensions and formulates the energy-scale ambiguity precisely. Section 4 derives the common $g/f$ rescaling for static horizons using both local-frame MDR and rainbow-metric viewpoints. Section 5 applies the result to the AC, MS, and G-DSR/GDRS cases and examines simple Schwarzschild implications, including the parametric size of the correction. Section 6 explains why agreement in the near-horizon temperature does not imply identical evaporation phenomenology, and Section 7 discusses the interpretation and limitations of the generalized DSR extension. We conclude with a summary of what has and has not been established.
@@ 2.3 Magueijo–Smolin (MS) DSR invariant
so that $g_{\rm MS}/f_{\rm MS}=1$. This equality will play an important role below: even though the MDR is non-trivial, the near-horizon temperature correction vanishes in the common $g/f$ scaling.
@@ 3.4 The central ambiguity: which energy enters $E_{\star}/E_{\mathrm{Pl}}$?
4 Static horizons and the common $g/f$ scaling of the Hawking temperature under a shared operational scale
@@ 6 Beyond the near-horizon temperature: sources of model dependence
The common scaling (23) at the surface-gravity/tunneling level does not imply that all DSR-motivated models lead to identical evaporation phenomenology. The Hawking temperature captures the near-horizon analyticity and redshift structure, but the observable flux at infinity depends on additional inputs. For a generic species $i$ with greybody factor $\Gamma_{i}(\omega)$, the standard energy flux has the schematic form
@@
Several concrete mechanisms can generate model dependence beyond the common near-horizon scaling. First, massive thresholds and the relation between energy, momentum, and group velocity can modify the emission rates even in cases where $T(E_{\star})$ itself is unchanged, such as models with $f=g$. Second, DSR constructions with curved momentum space can induce a non-trivial integration measure and therefore alter the density of states and the spectral distribution [26, 24]. Third, the greybody factors $\Gamma_{i}(\omega)$ can change if the effective wave equation, turning points, or propagation speed in the exterior region are modified; this effect can easily compete with or dominate a small temperature rescaling. Finally, in full DSR one expects non-linear composition laws for multi-particle kinematics, so the bookkeeping of emitted energy and the implementation of backreaction need not coincide with the standard semiclassical picture. These observations explain why agreement at the level of the near-horizon temperature does not imply agreement for integrated evaporation histories or for endpoint behavior.
The operational choice of $E_{\star}$ remains the dominant ambiguity even inside the temperature calculation itself. If $E_{\star}$ is identified with a detector energy at infinity, the deformation acts mainly as a frequency-dependent reweighting of the observed spectrum. If instead one ties $E_{\star}$ self-consistently to the local emission process, for example by taking $E_{\star}\sim T$, the correction becomes mass dependent and can qualitatively alter late-stage evaporation in models where $g/f$ approaches zero at finite argument [21, 20]. For the common estimate $E_{\star}=\xi T_{0}$ in the GDRS family, the relative correction is
$\frac{\delta T}{T_{0}}\simeq-(\Delta\alpha-\alpha_{2})\,\frac{\xi}{8\pi ME_{\mathrm{Pl}}}.$ (37)
Thus the effect is suppressed by the inverse black-hole mass in Planck units. As an order-of-magnitude illustration, one finds $|\delta T|/T_{0}\sim 4\times 10^{-40}|\Delta\alpha-\alpha_{2}|\,\xi$ for a solar-mass black hole, whereas even a much lighter black hole of mass $10^{15}\,$g still gives only $|\delta T|/T_{0}\sim 10^{-21}|\Delta\alpha-\alpha_{2}|\,\xi$. Any phenomenological bound derived from Hawking thermodynamics would therefore be weak unless one has access to primordial or genuinely Planckian black holes. This is precisely why a meaningful comparison across approaches must specify not only the prescription for $E_{\star}$ but also the dynamical and statistical ingredients beyond the near-horizon temperature.
@@ 7 Discussion: implications of the G-DSR/GDRS extension and interpretation of results
Our main technical result may be interpreted as follows. For static horizons, the temperature is governed by two ingredients: (i) the geometric pole structure in the radial momentum near the horizon, and (ii) the relation between energy and radial momentum implied by the MDR. Under the operational assumption that the deformation functions are evaluated at a fixed finite scale $E_{\star}$, ingredient (ii) enters only as an overall rescaling. Rainbow metrics implement precisely the same rescaling through an energy-dependent normalization of time and space in the orthonormal frame. Therefore, at the level of near-horizon temperature and under the shared-scale assumption, the two prescriptions are equivalent parametrizations.
@@ 8 Conclusion
We have compared two standard prescriptions for implementing DSR-motivated modified dispersion relations in black-hole spacetimes: (A) an energy-independent background with the MDR imposed in local orthonormal frames, and (B) the rainbow-metric construction in which the deformation is encoded in an explicitly energy-dependent effective geometry. The main objective of the paper has been to determine, in a precise and controlled way, whether these two prescriptions imply genuinely distinct near-horizon Hawking-temperature predictions or whether their apparent difference is mainly a matter of formal representation. For static, spherically symmetric horizons, we showed explicitly that both prescriptions lead to the same near-horizon modification of the Hawking temperature at the surface-gravity/tunneling level when expressed in the ffgg parametrization and evaluated at the same finite operational scale $E_{\star}$,
@@
Therefore, the main conclusion should be read as a conditional equivalence statement: once the physical prescription for $E_{\star}$ is fixed and shared, the distinction between “local-frame MDR on a fixed background” and “rainbow metric” becomes largely one of parametrization at the level of the near-horizon temperature.
@@
Just as importantly, the analysis makes clear what has not been established. The operational identification of the deformation scale $E_{\star}$ is not derived from first principles here; rather, it is an input that must be specified before the comparison can be carried out. Likewise, agreement in the near-horizon temperature does not fix the evaporation phenomenology, because thresholds, phase-space measures, greybody factors, and non-linear composition laws can all introduce additional model dependence. For the common choice $E_{\star}\sim T_{0}$, the relative correction is suppressed by $1/(ME_{\mathrm{Pl}})$ and is therefore tiny for macroscopic black holes, so any realistic phenomenological impact is expected only in primordial or nearly Planckian regimes where semiclassical control is itself limited.
Taken together, these results show that the present work should be viewed as a clarification of how two common curved-spacetime DSR implementations are related, together with a compact parametrization of the leading G-DSR/GDRS correction. Nevertheless, a fully predictive phenomenology will require going beyond the near-horizon temperature and deriving or motivating the relevant operational energy prescription in a more fundamental way.